40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Nov 13 22:32:54 2015, Koji, Summary, PSL, PMC LO degraded, usual ERA-5 replacement, LO recovered PMC_LO.pngIMG_2093.JPGIMG_2091.JPGIMG_2095.JPGIMG_2096.JPG
    Reply  Sun Nov 15 22:43:48 2015, Koji, Summary, PSL, PMC LO degraded, usual ERA-5 replacement, LO recovered C1-MULTI_E6875C_TIMESERIES-1131408017-86400.pngC1-MULTI_E6875C_TIMESERIES-1131580817-86400.png
       Reply  Mon Nov 16 19:05:59 2015, Koji, Summary, PSL, PMC servo circuit review, follow up measurements 
          Reply  Mon Nov 16 21:32:49 2015, Koji, Summary, PSL, PMC servo circuit review, follow up measurements 
             Reply  Tue Nov 17 16:21:10 2015, Koji, Summary, PSL, PMC servo circuit review, follow up measurements 
                Reply  Wed Nov 18 16:51:58 2015, Koji, Summary, PSL, PMC servo calibration PMC_err_cal.pdf
                   Reply  Wed Nov 18 16:53:17 2015, Koji, Summary, PSL, PMC PZT capacitance 
                      Reply  Wed Nov 18 17:09:22 2015, Koji, Summary, PSL, PMC Servo analysis PMC_OLTF.pdfPMC_DSP.pdf
                         Reply  Fri Nov 20 15:44:12 2015, Koji, Summary, PSL, Added 17.5kHz LPF to the PMC servo PMC_OLTF.pdfPMC_noise_comparison.pdf
Message ID: 11775     Entry time: Tue Nov 17 16:21:10 2015     In reply to: 11769     Reply to this: 11780
Author: Koji 
Type: Summary 
Category: PSL 
Subject: PMC servo circuit review, follow up measurements 

I'm still analyzing the open loop TF data. Here I report some nominal settings of the PMC servo

Nominal phase setting: 5.7
Nominal gain setting: 3dB

After the tuning of the notch frequency, I thought I could increase the gain from 5dB to 9dB.
However, after several hours of the modification, the PMC servo gradually started to have oscillation.
This seemed to be mitigated by reducing the gain down to 4dB. This may mean that the notch freq got drifted away
due to themperature rise in the module. PA85 produce significant amount of heat.

(The notch frequency did not change. Just the 22kHz peak was causing the oscillation.)

ELOG V3.1.3-